Should Adam Smith become a homeowner? Adam Smith completed his MBA at Eastern New Mexico University and accepted a job at Nissan North America Headquarter in Franklin, TN. Because Adam liked to have a very active social life, he decided to rent a two-bedroom condominium in downtown Nashville and commute to his job. The rent was $2,500 per month, which included parking but no utilities. Now, after one year, a similar condominium in his building has become available for sale with an asking price of $550,000. Adam likes the area, the building, and more important the unit itself, therefore he is considering buying the condominium. He knows he is not ready to settle down right now, but he is convinced that this will change in the next few years. Facing the classic buy vs rent dilemma, Adam has a long discussion with his best friend David Hume.

However, it is time for Adam to make his own decision, so he comes up with several financial details he must consider.

Adam believes he can purchase the condominium for $530,000. (A parking spot is included.)

The association fee will be $500 per month.

The property taxes will be $600 per month.

General maintenance and repairs will be $600 per year.

To purchase the unit, Adam will provide a 20% down payment.

The closing costs will be around $10,000.

Adam will finance the remaining 80% with a mortgage for 30 years.

The rate will be 4%.

The down payment is currently invested and brings a 4% annual return.

If he decides to sell the home, he will pay 6% of the selling price to the realtor plus $2,000 other fees. In order to perform a comprehensive analysis, Adam takes the following steps Compute the monthly mortgage payment. Compute the opportunity cost of using the required funds for closing. Compute the total monthly cost.

Perform a scenario analysis: compute net future gains if he sells the condominium in 2 years in 5 years 10 years Because he heard some rumors about Nashville’ house market being overpriced, Adam also considers a scenario analysis for future condo prices:

The price drops 10% over the next 2 years, reverts to the original price by the end of year 5, then increases 10% from the original price by the end of year 10.

The price increases by 5% annually.

Follow Adam’s steps and explain what he should do.

Did Adam miss anything?

What other factors can he consider?
Answer #1

I) Buy Vs. Rent Analysis:

Step 1: Monthly mortgage payment:

Loan amount = Purchase price - downpayment of 20%

= 530,000 - 20% (530,000)

= 530,000 - 106,000

= $424,000

Annual mortgage payment @ 4% rate = 424,000 * 4% = $16,960

Monthly mortgage payment = 16,960 /12 = $1,413.33

Step 2: Opportunity cost of using the funds required for closing:

Funds required for closing = down payment + closing costs

   = 106,000 + 10,000

Funds required for closing = $116,000

Opportunity cost = 4% annual return foregone by not investing the funds being used for closing

= 4% (116,000)

Opportunity cost per year = $4,640

Opportunity cost per month = 4,640/12 = $386.66

Step 3: Total monthly cost:

Total monthly cost = monthly mortgage payment + monthly association fees + monthly property taxes + monthly opportunity cost

= 1,413.33 + 500 + 600 + 386.66

= $2,513.33

So total monthly cost that will be incurred in case Adam buys the house = $2,513.33

The current rent cost =$2,500

Hence there is no significant change in the monthly cost if Adam buys the condominium instead of renting the curent one. He should instead consider the net future gains to decide on the buy vs. rent consideration, which again depends on the future scenario of the Nashville house market

II) Scenario Analysis for future Condo prices:

A) Sells condo in 2 years:

Price drops by 10% i.e. Sale price would be 90% (530,000) = $477,000

Future gain/(loss) = (Sale price - Purchase price) - 6 % realtor fees of selling price - other fees

   = (477,000-530,000) - 6% (477,000) - 2,000

   = (53,000) - 28,260 - 2,000

Future loss = $83,620

Hence in the scenario of selling the condo in 2 years, Adam will have a future loss of $83,620 and hence should not go for for the buy option in such a case

B) Sells condo in 5 years:

Price reverts to original price

However even this scenario will yield loss to Adam as he has to incur the realtor fees and other fees as shown below:

Future gain/(loss) = (Sale price - Purchase price) - 6 % realtor fees of selling price - other fees

   = (530,000-530,000) - 6% (530,000) - 2,000

   = 0 - 31,800 - 2,000

Future loss= $33,800

C) Sells condo in 10 years:

Price increases by 10% from the original price

So Sale price = 110% (530,000) = 583,000

Future gain/(loss) = (Sale price - Purchase price) - 6 % realtor fees of selling price - other fees

   = (583,000-530,000) - 6% (583,000) - 2,000

   = 53,000 - 34,980 - 2,000

Future gain = $16,020

Hence in this scenario, Adam should buy the condo as ir would fetch him future gains